A galvanic cell consists of a chromium, Cr bar in a chromic sulfate solution, Cr2(SO4)3 containing 13 g and a silver, Ag bar in a silver sulfate solution (Ag2SO4 ) containing 37.8 g in a

3 L solution and a sodium sulfate solution use as a salt bridge solution. The silver electrode is positive relative to the chromium electrode. Write the cell half- reactions, overall reaction and cell notation. Calculate the following:

A) Gibbs Free – Energy B) Electrode potential of the cell.

(Note: when calculating the molarity do not include the spectator ions)



1
Expert's answer
2021-12-07T08:51:02-0500

ΔG = ΔH − T * ΔS

Molar mass of Cr2(SO4)3=392.16

13/392.16=0.033mol

Molar mass of Ag2SO4=311.8

37.8/311.8=0.1212mol